OpenStudy (sawq):

This is an easy algebra 1 question!! will fan and medal!!!! Solve: 1/2x + 1 = –x + 8 i totally forgot how to do this??? please walk me through

OpenStudy (anonymous):

Well first you start off by trying to isolate the variable(x)

OpenStudy (sawq):

right. opposite operation?

OpenStudy (anonymous):

Right, To do this you can start by subtracting 1 from both sides to get \[1/2x =-x+7\]

OpenStudy (anonymous):

after that you have to subtract -x to both sides (add x) to get 3/2x=7

OpenStudy (anonymous):

then you divide both sides by 3/2 to get x = 4.6666 (repeating) or 14/3

OpenStudy (sawq):

but why 3/2?

OpenStudy (sawq):

@tbeneka

OpenStudy (anonymous):

it would be 3/2 because 1/2+1 = 3/2

OpenStudy (sawq):

but how does adding x get you 1?

OpenStudy (sawq):

@tbeneka

OpenStudy (sawq):

and also the options are: 7 1/2 14 18

OpenStudy (sawq):

@tbeneka pls help me finish sory im being annoying i just dont fully understand

OpenStudy (anonymous):

sorry my internet stopped working for a bit

OpenStudy (anonymous):

wait is the equation formatted like option one or two option one: \[\frac{ 1 }{ 2 }x+1=-x+8\] option two: \[\frac{ 1 }{ 2x }+1=-x+8\]

OpenStudy (sawq):

option 1

OpenStudy (anonymous):

Hmmm well I'm not sure because I'm a 100 percent sure the answer to number one is 14/3

OpenStudy (sawq):

ok thanks i chose 14 and got it right :)

OpenStudy (anonymous):

lol okay good job :DD

OpenStudy (sawq):

thanks for all ur help. gave you a medal and becam a fan